Grade 9Mechanics

The acceleration with which an object of mass 100 kg be lowered from roof using a cord with breaking strength of 60 kg weight without Breaking the rope is

surendra surendra
7 Years agoGrade 9

1 Answer

Arun
7 Years ago
Rope of the strength=mg=100×10=1000n
and breaking strength of rope is =mg=60×10=600
 
Hope it helps
 
Regards
Arun